OpenStudy (anonymous):

James left a bin outside in his garden to collect rainwater. He notices that 1 over 4 gallon of water fills 2 over 3 of the bin. Write and solve an expression to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

OpenStudy (adamaero):

1/4 gallon fills 2/3 bin

OpenStudy (adamaero):

.25x = 2/3 3*(.25)x = 2 .75x = 2 3/4x = 2 (divide each side by 2) 3/8x = 1 x = 8/3 (gallons) If that makes sense, then it may be right. Although, I don't care that much (tired) so there's a high chance this answer is wrong.